# Dynamic Recourse Allocation with Karma: An Experimental Study

### v2

@@ Abstract
journal: Games and Economic Behavior
@@
[uzh] organization=Department of Sociology, University of Zürich,addressline=Andreastrasse 15, city=Zürich, postcode=8050, country=Switzerland
@@ 1 Introduction
Figure 1: Dynamic resource allocation with karma involving three cars and a repeated sequence of three encounters. Start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 6, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. The bids of the lila car (and, in general, of other cars winning in parallel encounters) are paid and redistributed. As a result, the blue car’s karma account goes up by 3 to 12. Let’s move along clockwise. Blue now happens to have high urgency and bids 6, thus outbidding and getting priority over orange whose karma account is 9 and bids 3. Now the orange karma goes up by the redistribution share of 3 to 12. Let’s move along clockwise. Orange now has high urgency and bids 6, thus outbidding low-urgency lila who has 6 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
@@
The choice system, which has indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations (Prendergast, 2022), provides empirical evidence that karma mechanisms can be successful in practice. On the theory side, recent works have studied game-theoretical models of karma mechanisms in simplified, single repeated resource allocation settings (Elokda et al., 2023, 2024). These works have been motivated by the use of karma as a public policy instrument that is an equitable alternative to classical (monetary) congestion pricing policies, e.g., for allocating priority roads and other public infrastructures. However, in order to realize the potential benefits of karma-based policies, and employ these policies effectively in human populations, systematic behavioral evidence is needed. The present paper contributes to this emergent strand of literature a first controlled experimental test of karma. We conduct an experiment on karma with different bidding schemes and different urgency processes in order to better understand the behavioral uptake of karma by human actors and its efficiency and distributional consequences.
@@
The theoretical predictions of Elokda et al. (2023, 2024) are based on the solution concept of the Stationary Nash Equilibrium (SNE): a compact, time-invariant predictor of optimal rational behavior, which is guaranteed to exist due to the preservation of karma. Elokda et al. (2023) develop computation tools for the SNE and show that farsighted SNE are almost fully efficient with respect to the private urgency of the users. These predictions are based on idealized assumptions including rationality, far-sightedness, infinite population, perfect adoption, and that the stationary conditions for which the SNE is optimal are reached. In this paper, we perform a behavioral investigation with real humans who do not necessarily follow such idealized assumptions. We want to know whether humans find it natural to adopt the mechanism, whether inexperienced players are able to realize efficiency gains, and whether stationarity is attained in practice. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic urgency distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as a benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private urgency, including fixed turn-taking schemes and schemes using fixed-value tokens.
@@
    Almost all participants benefit in the karma allocation compared with random allocation: this is the case for 90% of the participants, and the remaining 10% are mostly non-adopters who do not participate in karma bidding actively themselves. If we look at active participants only, karma led to an almost Pareto improvement.
@@
    The Pareto improvement occurs despite the relatively low level of training and commitment by the online participants.
@@
    The realized benefits, while greater than in random allocation, fall short of theoretical Nash predictions. Analysis of the bidding behaviors reveals that the main deviation to Nash behavior takes the form of irrational over-bidding in low urgency rounds.
@@
    Benefits are particularly pronounced in situations when preference intensities are dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary).
@@
These findings provide a first benchmark that karma may be used beneficially and robustly in human interactions. Our study also points in several directions for further theoretical and experimental investigation.
@@ 2 Related Mechanisms
Mechanisms based on linking decisions (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013) rely on correlating each individual’s reports over time with publicly known distributions of the private urgency, and punishing those individuals that deviate. This requires to keep track of individual identities and histories, and we view karma and other tokens as memory-efficient instruments to link decisions that are capable of scaling in large populations.
@@
Trading votes across issues or proposals is an intuitively appealing and practically prevalent practice, yet it remains unclear to what extent vote trading improves welfare and how to design vote trading mechanisms optimally, as pointed out in Casella and Macé (2021)’s recent review. Casella and Macé (2021) distinguish between two types of vote trading: those in which votes are traded with other voters (Casella and Palfrey, 2019, 2021); and those in which votes are traded individually with one’s future self (referred to as storable (Casella, 2005) or qualitative votes (Hortala-Vallve, 2012)). The latter type of storable votes, which yields particularly favorable efficiency gains in comparison to the other types of vote trading (Casella and Macé, 2021), is closely related to the aforementioned class of (finite-horizon) artificial currency mechanisms: voters are issued an initial budget of votes to cast in a (small) finite number of issues (Casella, 2005; Casella et al., 2006; Hortala-Vallve, 2012). One recently proposed mechanism by Macías (2024) resembles karma more closely: in this mechanism, votes are “paid” by the majority voters and subsequently redistributed, however, Macías (2024) studies a two voter only model. Our study thus complements the literature on storable votes, as we are motivated by resource allocations that are repeated more frequently and typically involve more players than in voting. In contrast to previous experimental studies on vote trading (Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Casella and Palfrey, 2019), our experiments involve significantly more rounds and larger groups.
@@ 3 Experimental Methods
3.1 The Game
The game we study is one that proceeds with $N$ participants over $T$-many rounds. All participants receive an initial endowment of karma $k(0)=k(1)={k_{\textup{init}}}\in\mathbb{N}$ and an initial game score $s(0)=0$. At each round $t\in\{1,\dots,T\}$, participants are randomly matched in pairs to compete over a shared resource. Each participant is given an urgency value $u(t)\in\{{u_{\textup{l}}},\>{u_{\textup{h}}}\}$ that is drawn randomly and independently from a process $\mathbb{P}(u)$ that is identical for all players, and must place a bid $b(t)\leq k(t)$. In each pairwise matching, the higher bidder gets allocated the resource: the score $s(t)=s(t-1)+u(t)$ increases by the urgency and the bid is collected as payment. The lower bidder does not get allocated the resource: the score $s(t)=s(t-1)$ does not increase and no payment is collected. Ties are settled randomly. To keep the total amount of karma constant, at the end of each round, the total collected payment ${p_{\textup{tot}}}$ is uniformly redistributed to all participants in an integer preserving manner. In case ${p_{\textup{tot}}}$ is not divisible by $N$, $\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ is redistributed to everyone, and a random subset of ${p_{\textup{tot}}}-N\left\lfloor\frac{{p_{\textup{tot}}}}{N}\right\rfloor$ participants receive one additional karma. Moreover, each participant is allowed a maximum karma level ${k_{\textup{max}}}$ that the redistribution respects. Participants with ${k_{\textup{max}}}$ do not receive additional redistribution which instead gets issued uniformly to the others. After the redistribution the game proceeds to the next round $t+1$.
Table 3 in A summarizes the notation introduced above.
3.2 Treatments
We follow a two-by-two factorial treatment design, where we vary the dynamic urgency process of the participants and the richness of the karma scheme. Table 1 summarizes the treatment configurations schematically. For each of the four resulting treatments we run 5 independent game experiments with 20 participants per experimental game resulting in 100 subjects per treatment and 400 subjects in total.
Table 1: 2x2 experimental design.
Richness of scheme
Binary Full range
Urgency process
3.2.1 Urgency Process
3.2.2 Richness of Karma Scheme
In the treatment variation of binary, participants can only choose between two bid levels: 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ (i.e., half their karma, rounded down to the nearest integer). In the treatment variation of full range, participants can choose any integer bid up to their karma. The motivation of these treatment variations is to investigate the behavioral effects of a reduced action space, i.e., binary, where subjects either bid or not. The particular design choice to restrict bids to 00 or $\left\lfloor\frac{k}{2}\right\rfloor$ in the binary scheme was guided by theoretical Nash predictions: using tools from Elokda et al. (2023), these binary bid levels were predicted to achieve almost the same efficiency at the Nash equilibrium as in full range bidding, for the urgency processes considered (cf. Nash efficiency gains in Figure 2). Therefore, the simpler binary scheme does not trade-off performance, in theory, and any differences observed in the outcomes of the two bidding schemes will be due to behavioral effects.
3.3 Experimental Implementation
The game is implemented as a real-time online experiment using oTree (Chen et al., 2016). Participants are given the opportunity to familiarize with the game over $5$ rounds that do no contribute to the game score, after which the main game proceeds with $T=50$ rounds. In each round, participants get drawn a random urgency. On the decision page, they must place a bid, and then a results page follows which provides feedback on the outcomes of that round. The bid must be placed within $10$ seconds, otherwise the participant is signalled as inactive, and the bid defaults to zero. The results page communicates the outcome of the round in terms of whether priority is granted, the karma payment and redistribution, and the updated karma balance and game score. In addition, it gives feedback on the opposing bid (but not the opposing urgency and karma). Figures 78 in A show examples of these pages.
At the end of the game, participants are awarded a monetary payoff consisting of a fixed fee ${\phi_{\textup{fix}}}$ and a bonus fee ${\phi_{\textup{bon}}}$ that depends on the final score $s(T)$. Therefore, they are encouraged to be the winning bidder in as many rounds as possible, and especially in high urgency rounds. The bonus fee is determined according to the following rule:
  • 1.
    If the participant is inactive for more than $6$ consecutive rounds, they are considered to not complete the experiment and are not awarded any payoff, i.e., ${\phi_{\textup{fix}}}={\phi_{\textup{bon}}}=0$;
  • 2.
    Otherwise, the bonus fee is computed as an affine function of the score, given by
This rule linearly interpolates between two payment levels: ${\phi_{\textup{targ}}}$ is the payment associated with a target score ${s_{\textup{targ}}}$; and ${\phi_{\textup{rand}}}$ is the payment associated with the expected score for random bidding ${s_{\textup{rand}}}$. Hence, setting ${\phi_{\textup{rand}}}$ to a low value disincentivizes random play.
Participants were recruited on Amazon Mechanical Turk (MTurk). For each of the four treatment combinations, we ran five experiments with $N=20$ participants each, for a total of 100 participants per treatment and 400 participants overall. The fixed fee was ${\phi_{\textup{fix}}}=\$1.5$ and compensates for a maximum waiting time of 10 minutes to form an experiment group. The bonus fee ${\phi_{\textup{bon}}}$ was tuned based on the observed performance in technical pre-tests such that participants receive approximately ${\phi_{\textup{targ}}}=\$10$ on average, which compensates for a maximum experiment duration of 40 minutes.
Table 3 in A lists all parameter values used in the experiments.
4 Results
In order to present our findings, we must first introduce the welfare measures used to quantitatively assess our results.
4.1 Welfare Measures
Our central welfare measure is the efficiency gain which we define next. For a particular participant $i$, let $\left(u_{i}(t)\right)_{t\in\{1,\dots,T\}}$ be the vector of realized urgency in the experiment, $S_{i}=s_{i}(T)$ the total score at the end of the experiment, and $S^{\textup{rand}}_{i}=\frac{1}{2}\sum_{t=1}^{T}u_{i}(t)$ the expected total score, given the urgency realization, under random allocation. Then the efficiency gain of participant $i$ is defined as
$E_{i}=\frac{S_{i}-S^{\textup{rand}}_{i}}{S^{\textup{rand}}_{i}}.$ (1)
This definition expresses the relative improvement with respect to the expected random score given the urgency realization, in order to control for randomness in the urgency realization. On this basis, we will assess overall efficiency based on the median efficiency gain among participants, and fairness based on the distribution of efficiency gains.
[…]
4.2 High-level Synopsis of Results
Overall, we find that there are pronounced and statistically significant efficiency gains in all treatments, higher under the high stake process than under the low stake process, but not significantly different from one another in terms of overall efficiency gains depending on whether the bidding scheme is binary or full range. In all treatments, more than $90\%$ of the population is better off with karma than under random allocation. The most favorable combination is high stake with binary bidding, both in terms of median efficiency gains and in terms of the distribution of gains as most individuals achieve pronounced benefits.
Remark: With the best intentions, we had fully pre-registered design and analysis of our experiments. Some but not all of the analyses presented here were indeed pre-registered, and there are also further analyses from the pre-registration that are not presented in this paper. We stuck with the pre-registration as much as we could, but the unfortunate need to depart from the pre-registered plan of analysis is explained in B.
4.3 Efficiency Results
Efficiency went up in all treatments. Figure 2 shows the median efficiency gain in each of the four treatments (denoted “Total”), cf. Table 1, as well the median efficiency gain in the bottom and top halves of the population in terms of individual outcomes, contrasted to the median gains that would be realized if all participants followed theoretically optimal Nash behavior. For the purpose of visualizing the data spread, estimates of the $95\%$ confidence intervals (CI) are also shown in the figure based on 1000 bootstraps [^3] Note that data samples are dependent for participants in the same experiment, therefore the bootstrapped CIs should be seen as heuristic estimates for visualization purposes. Similar estimates were attained for larger numbers of bootstraps.. For the random and Nash allocations, the CI was estimated by running 1000 independent simulations.
The findings of Figure 2 summarize as follows. Although the experimental efficiency gains fall short of the theoretical Nash gains, there are nonetheless positive total efficiency gains in all treatments (median ranging from $7.38\%$ to $15.34\%$). The efficiency gains are especially pronounced in the upper half of the population (median ranging from $22.51\%$ to $29.85\%$); whereas the lower half of the population does not perform much worse than ex-ante expected under random allocation (median ranging from $-6.5\%$ to $-0.65\%$), and in fact performs better than the lower half of the population under random allocation ex-post. This suggests that overall, karma provides an opportunity for participants to achieve pronounced benefits, without strongly harming those that are less strategic. Moreover, the efficiency gains are statistically significant in all treatments, as elaborated in Table 2 reporting MWW test statistics $U$ (with associated $p$-values).
Figure 2: Median efficiency gains for the four treatment combinations, cf. Table 1, with estimate of $95\%$ confidence interval. Also shown are the medians in the lower and upper halves of the population.
Table 2: Mann–Whitney–Wilcoxon (MWW) tests with test statistics $U$ (and associated $p$-values in brackets) of efficiency gains under karma versus random for all treatments. Karma treatments have $n_{\textup{karma}}=100$, and random is based on $n_{\textup{random}}=100\,000$ samples.
Richness of scheme
Binary Full range Combined
Urgency process
Low stake
6 187 810.06187810.06\,187\,810.06 187 810.0
($<0.001$)
6 591 766.56591766.56\,591\,766.56 591 766.5
($<0.001$)
25 557 373.025557373.025\,557\,373.025 557 373.0
($<0.001$)
High stake
6 640 154.56640154.56\,640\,154.56 640 154.5
($<0.001$)
6 242 150.06242150.06\,242\,150.06 242 150.0
($<0.001$)
25 765 437.025765437.025\,765\,437.025 765 437.0
($<0.001$)
Combined
25 701 703.025701703.025\,701\,703.025 701 703.0
($<0.001$)
25 566 897.025566897.025\,566\,897.025 566 897.0
($<0.001$)
102 525 023.5102525023.5102\,525\,023.5102 525 023.5
($<0.001$)
[…]
4.4 Fairness Results
Figure 2 shows that in addition to achieving overall efficiency gains, the karma allocation is more efficient than random allocation for both the lower/less fortunate and the upper/more fortunate halves of the population. To provide finer grained insight on the fairness of karma, Figure 3 further shows the mean efficiency gain per population decile for the four treatments. It is important to note that while the efficiency gain (1) is defined with respect to the ex-ante expected score under random allocation given the urgency realization, not the whole population realizes this score ex-post. To control for the fact that there will be more or less fortunate individuals due to radnomness under random allocation, Figure 3 also shows the ex-post mean efficiency gains per population decile for random allocation, with the $95\%$ confidence interval estimated from 1000 independent simulations per treatment. The key feature to observe is that in all treatments, $90\%$ of the population achieve higher efficiency in the karma scheme than the random scheme. Only the lowest decile is worse-off in karma than random. This decile is dominated by non-adopters whose bids defaulted to zero due to inactivity (number of non-adopters: low stake-binary: 9/100; low stake-full range: 6/100; high stake-binary: 7/100; high stake-full range: 5/100). Non-adopters achieve particularly low scores as with a consistent bid of zero it is very unlikely to get granted priority.
Figure 3: Mean efficiency gain per decile for the four treatment combinations.
Moreover, it is evident that the treatment combination of high stake-binary is particularly favorable. Not only does it lead to the highest total median efficiency gain, cf. Figure 2, but it also achieves the largest gap to random allocation across deciles, cf. right panel of Figure 3. This suggests that all (active) subgroups of the population achieve pronounced benefits in this treatment.
Notice that the variability in the ex-post efficiency gains under random allocation, cf. Figure 3, can be improved by adopting simple turn-taking schemes, such as a fixed-value token scheme granting priority to participants holding the highest number of tokens (or equivalently, the lowest number of previous allocations). Such a scheme would achieve the same aggregate efficiency as random allocation, but ensure the ex-post efficiency gains are more equally distributed around zero in Figure 3. Therefore, in comparison to a simple turn-taking scheme, karma would also achieve pronounced efficiency gains overall, meanwhile the proportion of the population benefiting would drop to $70\%$ to $80\%$ depending on the treatment.
4.5 Analysis of Bidding Behaviors
[…]
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 4: Median bid per urgency and karma for the four treatment combinations.
The main finding of Figure 4 is that overall, there was a tendency to over-bid in the low urgency state [^4] Notice that due to the aforementioned bug which excluded some of the losing bids, the medians in Figure 4 are potentially biased towards higher winning bids. Nonetheless, there are sufficiently many samples to support the observed tendency to over-bid in low urgency.. The bidding behaviors are not well-explained by a single value of future discounting: the median bids are well-fitted to far-sighted behavior (discount factors 0.80.80.80.80.980.980.980.98) in high urgency, but to short-sighted behavior (discount factor $0.6$) in low urgency. In contrast, the median bids of top-performing participants are well-fitted to far-sighted behavior in both urgency levels, which explains the high efficiency gains achieved by these participants.
To provide further insight on how over/under-bidding during low/high urgency affected performance, Figure 5 shows a scatter of the mean signed difference to Nash-$0.98$ bids per participant and urgency level, versus the attained efficiency gains, for the four treatments. The data points associated to the top-performing decile are highlighted in orange. As expected, under-bidding in high urgency negatively affects performance, cf. third quadrant of the high urgency panels, which includes the majority of participants achieving negative efficiency gains. On the other hand, the effect of over-bidding in either of the urgency levels is inconclusive. Consistently in all treatments, the single-top performing participant bid close to Nash [^5] The mean absolute difference to Nash-$0.98$ bids correlates well with the efficiency gain, as expected, and particularly so in the full-range treatments (Spearman correlations: low stake, binary: $-0.345$, $p=0.0004$; low stake, full-range: $-0.664$, $p=0.0000$; high stake, binary: $-0.119$, $p=0.236$; high stake, full-range: $-0.674$, $p=0.0000$). However, several top performing participants also tended to over-bid in low urgency, and a few of these participants managed to over-bid in both low and high urgency, cf. Figure 5(d).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 5: Efficiency gain vs. mean difference to Nash for the four treatment combinations.
4.6 Analysis of stationarity
An important feature of the karma mechanism is that it forms a closed economy in which karma is preserved over time. In theory, this feature enables reaching a stationary regime, in which a predictable, time-invariant optimal behavior can be repeated indefinitely. To investigate whether stationarity is achieved in practice, each sub-figure of Figure 6 shows the variations in the distribution of karma (left panel) and bids (right panel) over the course of the main experimental rounds, for each of the four treatments. The variation in the distributions is measured by the Wasserstein-1 distance, also known as the Earth mover’s distance, between the realized karma/bid distributions of successive rounds. The choice of the Wasserstein-1 distance is interpretable: for example, a distance of one corresponds to shifting the distribution by one karma unit on average. For each treatment, the realized distribution variations are plotted for each of the five experimental groups (labelled “MTurk-G1–G5”), and contrasted to the variations that would be attained if all participants followed stationary Nash behavior (labelled “Nash”; the shaded area coincides with the 95% confidence interval estimated from 1000 independent simulations per group).
(a) Low stake, Binary.
(b) Low stake, Full range.
(c) High stake, Binary.
(d) High stake, Full range.
Figure 6: Variation in the karma and bid distributions over time for the four treatment combinations.
The main finding of Figure 6 is that in all treatment combinations, after a short initial transient of less than five rounds, the realized variations in the karma and bid distributions are close in magnitude to those attained under stationary Nash play, as visualized by the shaded Nash area subsuming most of the experimental data. Two exceptions are the bid distributions in the full range treatments, cf. right panels of Figures 6(b), 6(d), in which the experimental variations lie at a notably higher magnitude than the stationary Nash variations. Nonetheless, these variations are still relatively small and oscillate around a Wasserstein-1 distance of one. Overall, this finding suggests that despite the presence of noisy bids, the karma auctions quickly become predictable over time, and these auctions are even more predictable in the binary treatments.
5 Discussion
In sum, in this paper we find that the aggregate efficiency gains of a karma scheme compared with random allocation are pronounced and statistically significant in all treatments. This constitutes the first set of behavioral evidence that a formal karma mechanism indeed can work to the benefit of the population. Moreover, in all treatments, (almost) all participants manage to benefit from the karma scheme in comparison to random allocation, with the exception of the lowest decile of non-adopters who did not actively participate in the bidding. Thus, our experiments provide the first set of formal evidence for the potential social benefits of using a karma scheme with human participants, as it improves efficiency to the benefit of almost everybody.
It is noteworthy that our experimental subjects were recruited from a population of totally untrained and inexperienced users from an online convenience sample (on MTurk). Their behavior, even though more efficient than random, however, is not as efficient as is theoretically feasible under Nash equilibrium play. Analysis of the experimental bidding behaviors reveals a consistent tendency to over-bid in low urgency rounds, however, this observation does not explain the gap to Nash efficiency entirely, which is likely attributed to irrational, noisy behavior. A natural follow-up question is, therefore, whether karma could be capable of achieving higher efficiency gains than realized in our online experiments if the human population consisted of participants that were better trained. As a first step in this direction, we conducted an auxiliary experiment under the low stake-full range treatment with a group of ‘expert’ subjects, who were graduate students in an applied game theory class. It turned out that, indeed, the achieved efficiency gains of these experts are close to Nash levels (MTurk: median $12.86\%$, $n=100$; experts: median $36.65\%$, $n=28$, MWW test $U=667.0$, $p=0.0000$). Therefore, we may interpret the efficiency gains that were realized in our online experiment as behaviorally robust lower bounds on the performance of the karma scheme given the relatively low training and commitment of the subject pool considered.
Finally, to conclude, we would like to highlight that the most favorable treatment combination in our experiments was that of high stake urgency process under binary bidding scheme; both median efficiency gains and the distribution of gains were higher in that combination than in all others. Having rare but important urgency realizations made it particularly easy for our subjects to decide whether to bid or not, and there were no subtleties regarding how much to bid as a function of the history of play, etc. given the binary nature of the bidding scheme.
Appendix A Supplementary Implementation Details
Figure 7 shows examples of the decision page implemented in the online experiment. Figure 8 shows an example of the results page. Table 3 lists the notation and detailed parameter values used in the experiments.
(a) High stake, binary treatment.
(b) Low stake, full range treatment.
Figure 7: Decision page examples.
Figure 8: Results page example.
@@ Appendix A Supplementary Implementation Details
${T_{\textup{dec}}}$ Decision inactivity timer 10s

### v3

# Dynamic Resource Allocation with Karma: An Experimental Study

@@ Abstract
journal: Journal of Economic Behavior & Organization
@@
[uzh] organization=Zurich Center for Market Design and SUZ, University of Zürich,addressline=Andreastrasse 15, city=Zürich, postcode=8050, country=Switzerland
@@ 1 Introduction
Figure 1: Dynamic resource allocation with karma involving three cars and a repeated sequence of three encounters. Start with the encounter in the top center. The high-urgency lila car has a current karma account of 9 and bids 6, thus outbidding the low-urgency blue car whose karma account is also 9 but bids 2. The bids of the lila car (and, in general, of other cars winning in parallel encounters) are paid and redistributed in the population. As a result, the blue car’s karma account goes up by 3 to 12. Let’s move along clockwise. Blue now happens to have high urgency and bids 6, thus outbidding and getting priority over orange whose karma account is 9 and bids 3. Now the orange karma goes up by the redistribution share of 3 to 12. Let’s move along clockwise. Orange now has high urgency and bids 6, thus outbidding low-urgency lila who has 6 karma left and bids 1. Thus the circle closes, et cetera, et cetera.
@@
The choice system, which has indeed resulted in significant aggregate gains in terms of food bank participation and fluidity of donations (Prendergast, 2022), provides empirical evidence that karma mechanisms can be successful in practice. On the theory side, recent works have proposed game-theoretic models of karma mechanisms in repeated resource allocation settings (Elokda et al., 2023, 2024b). These works have been motivated by the use of karma as a public policy instrument that is an equitable alternative to classical (monetary) congestion pricing policies, e.g., for allocating priority roads and other public infrastructures. However, in order to realize the potential benefits of karma-based policies, and employ these policies effectively in human populations, systematic behavioral evidence is needed. The present paper contributes a first controlled experimental test of karma to this emergent strand of literature. We conduct an experiment on karma with different bidding schemes and different urgency processes in order to better understand the behavioral uptake of karma by human actors and its efficiency and distributional consequences.
@@
The theoretical predictions of Elokda et al. (2023, 2024b) are based on the solution concept of the Stationary Nash Equilibrium (SNE): a compact, time-invariant predictor of optimal rational behavior, which is guaranteed to exist due to the preservation of karma. Elokda et al. (2023) develop computation tools for the SNE and show that farsighted SNE are almost fully efficient with respect to the private urgency of the users. It focuses on the setting of repeated pairwise matches competing over a single resource, which generalizes the stylized intersection example illustrated in Figure 1 for a continuum of agents from which pairwise matches are randomly drawn. We note, however, that the same theoretical and computational framework is applicable to more general settings in which more than two agents compete over more than one resource in each time-step (Elokda et al., 2024b, c). The theoretical predictions in these works are based on idealized assumptions including rationality, far-sightedness, perfect adoption, infinite population, and that the stationary conditions for which the SNE is optimal are reached. In this paper, we perform a behavioral investigation with real humans who do not necessarily follow such idealized assumptions. In particular, we recruit subjects from an online sample via Amazon Mechanical Turk (MTurk), sometimes criticized for more randomness in the resulting choice data, which for us is not a concern as we are interested in evidence that is robust to noisy and irrational behaviors of non-experts in such situations. We adopt the randomized pairwise matching setting of Elokda et al. (2023) to focus our investigation on the fundamental decision problem of trading-off present versus future resource consumption with karma. We want to know whether the inexperienced online subjects find it natural to adopt the mechanism, are able to realize efficiency gains, and whether stationarity is attained in practice. In our analysis we focus on the overall efficiency effects of karma as well as on distributional consequences and associated fairness properties of the mechanism. In our treatments, we vary the nature of the dynamic urgency distributions (more frequent and less intense versus less frequent and more intense) and the auction process of the karma mechanism (binary bidding versus full bidding). We use random allocation as the main benchmark to compare our findings to. This domain-independent benchmark is representative of common schemes that are unaware of the individual private urgency, including fixed turn-taking schemes and schemes using fixed-value tokens.
@@
    Almost all participants benefit in the karma allocation compared with random allocation: this is the case for 90% of the participants, and the remaining 10% are mostly dropouts who do not participate in karma bidding actively themselves. If we look at active participants only, karma led to an almost Pareto improvement.
@@
    The Pareto improvement occurs despite the relatively low level of training and commitment by the online participants and the variations in strategic competencies observed for different participants.
@@
    The realized aggregate benefits, while greater than in random allocation, fall short of theoretical Nash predictions. Analysis of the bidding behaviors reveals that close to Nash-level benefits are attained by participants that deviated from Nash by one karma bid unit on average, while positive benefits are still attained by participants that deviated from Nash by up to 3–4 karma bid units on average. Moreover, the main deviation to Nash behavior takes the form of irrational over-bidding in low urgency rounds.
@@
    Benefits are particularly pronounced in situations when high urgency is dynamically more intense and less frequent, and the bidding scheme is designed to be minimal (i.e., binary).
@@
These findings provide a first formal evidence that karma may be used beneficially and robustly in human interactions. Our study also points in several directions for further theoretical investigation and serves as a benchmark for future experiments.
@@ 2 Related Mechanisms
Mechanisms based on linking decisions (Jackson and Sonnenschein, 2007; Hortala-Vallve, 2010; Escobar and Toikka, 2013) rely on correlating each individual’s reports over time with publicly known distributions of the private urgency, and punishing those individuals that deviate. A variant of these linking mechanisms was tested behaviorally by Engelmann and Grimm (2012), which reports near-optimal efficiency gains by the experimental subjects as theoretically predicted. The mechanism in Engelmann and Grimm (2012) uses exact knowledge of the private urgency process and issues a tailored amount of fixed-value tokens, i.e., for the considered setting of 50-50% chance of low or high urgency the mechanism issues 20 tokens to spend over 40 periods. In contrast, karma mechanisms are able to adapt to any private urgency process, through a bidding scheme that does not fix the value of karma. In our experiments, we subject the same karma mechanism to different treatments varying the private urgency processes.
@@
Trading votes across issues or proposals is an intuitively appealing and practically prevalent practice, yet it remains unclear to what extent vote trading improves welfare and how to design vote trading mechanisms optimally, as pointed out in Casella and Macé (2021)’s recent review. Casella and Macé (2021) distinguish between two types of vote trading: those in which votes are traded with other voters (Casella and Palfrey, 2019, 2021); and those in which votes are traded individually with one’s future self (referred to as storable (Casella, 2005) or qualitative votes (Hortala-Vallve, 2012)). The latter type of storable votes, which yields particularly favorable efficiency gains in comparison to the other types of vote trading (Casella and Macé, 2021), is closely related to the aforementioned class of (finite-horizon) artificial currency mechanisms: voters are issued an initial budget of votes to cast in a (small) finite number of issues (Casella, 2005; Casella et al., 2006; Hortala-Vallve, 2012). One recently proposed mechanism by Macías (2024) resembles karma more closely: in this mechanism, votes are ‘paid’ by the majority voters and subsequently redistributed, however, Macías (2024) studies a two voter only model. Our study thus complements the literature on storable votes, as we are motivated by resource allocations that are repeated more frequently and typically involve more players than in voting. In contrast to previous experimental studies on vote trading (Casella et al., 2006; Hortala-Vallve and Llorente-Saguer, 2010; Casella and Palfrey, 2019), our experiments involve significantly more rounds and larger groups.
Summary of related mechanisms
To summarize, the present state of literature does not offer any natural mechanisms that simultaneously a) can be repeated indefinitely; b) are able to elicit private time-varying preferences; and c) do not require exact public knowledge of potentially private preference dynamics to do so. Karma mechanisms fill this gap. The lack of natural alternatives furthermore motivates our choice of adopting random allocation as the main benchmark in this first behavioral investigation of karma.
@@ 3 Experimental Methods
$T_{\textup{dec}}$} Decision inactivity timer 10s (test rounds: 20s)
$T_{\textup{res}}$} Results inactivity timer 15s (test rounds: 20s)